OpenStudy (anonymous):

Which expression is equivalent to the following complex fraction? mc019-1.jpg

OpenStudy (anonymous):

\[1+\frac{ 1 }{ y }\div 1- \frac{ 1 }{ y }\]

OpenStudy (anonymous):

@Mathhelp123344

OpenStudy (anonymous):

@mathmale

OpenStudy (anonymous):

@IMstuck

OpenStudy (anonymous):

@cwrw238

OpenStudy (mathmale):

\[1+\frac{ 1 }{ y }\div 1- \frac{ 1 }{ y }\]...if interpreted exactly as you have typed it, would require dividing (1/y) by 1, because we MUST perform multiplication and division BEFORE addition and subtraction. Dividing (1/y) by 1 doesn't make much sense, does it? I suspect that what you actually have here is\[\frac{ 1+\frac{ 1 }{ y } }{ 1-\frac{ 1 }{ y } },\]

OpenStudy (mathmale):

.../which could also be expressed, using parentheses liberally, as (1 + 1/y) / (1 - 1/y). In any case, if your readers have to figure out your possible meanings, valuable time will be lost. So I'm asking you to go back to the original problem and to type it again here, using parentheses or Equation Editor, in such a way that leaves no doubt regarding your meaning. Then we can discuss how to solve the problem.

OpenStudy (anonymous):

What is the quotient of \[\frac{ 2^{4} }{ 2^{-4}}\]

OpenStudy (anonymous):

@mathmale

OpenStudy (anonymous):

would the answer be 256

OpenStudy (mathmale):

Would you mind showing or explaining how you obtained that '256'?

OpenStudy (mathmale):

Hint:\[\frac{ 2^a }{ 2^b }=2^{a-b}\]

OpenStudy (anonymous):

2^4 = 16 2^-4 =.0625 and i just divided both those answers

OpenStudy (anonymous):

OHhhhhh ok ive been doing this wrong all along

OpenStudy (anonymous):

so the answer would be 1

OpenStudy (mathmale):

that works fine and is legitimate. If you use the example I gave you (above), you'll get\[\frac{ 2^4 }{ 2^{-4} }=2^{4-(-4)}=2^8=??\]

OpenStudy (mathmale):

No, it's not 1. Try finishing the calculation above. what is 2^8?

OpenStudy (anonymous):

so 256

OpenStudy (mathmale):

Right. 2^8 = 256. 2^8 also equals (2^4)^2=16^2 = 256. You OK with this?

OpenStudy (anonymous):

Which value is equivalent to \[\frac{ x+4 }{ 4+x }\]? –4 –1 1 4

OpenStudy (anonymous):

yes i am

OpenStudy (anonymous):

@mathmale

OpenStudy (mathmale):

Regarding:\[\frac{ x+4 }{ 4+x }\]...I'd prefer you post each new question in the "Ask a question" box. But anyway, ...\[\frac{ x+4 }{ 4+x }\] can be re-written as \[\frac{ x+4 }{ x+4 }\]...without changing its value. Please reduce this.

OpenStudy (anonymous):

wouldnt it just be 1